please help me need to check homework to see if correct working out

21 o.ZH 8,001 2 CI 2 S2
Answer #1

12SL 100 Rad/sec. 0.2H 0. 001 Impedonce of parollel RLC circuit 2 100X 0.2. = 20 Ω = 1-10.309.2n complezfoxm 20 -11o 20 0 12 x -j20 12-。20 K4 58. 823 -5.294 = 10.289-30.962.1lt) Vt) /ZI 75WSし100t) 10-289 1b) tlt)